OpenStudy (anonymous):

Hello,Can someone help me, I need to help my 9 year old solve this problem. Add and subtract. write in simplest form 3\8+5\6 ,and 11\-3\4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

3\8+5\6 = n [8] 3 +40/6 = 8n [6] 18 + 40 = 48n 58 = 48 n 58/48 = n = 29/24

OpenStudy (amistre64):

11/-3/4 11/-12 = -11/12